OpenStudy (drewbooks):

How do you simplify this? x^2yz/y^2 * y/2x

OpenStudy (anonymous):

I used a online calculator and got x^3z all over 2. I am pretty sure that this is correct but someone might need to check my answer.

OpenStudy (drewbooks):

That may be a possible solution, but it's not one of the options. Here they are:

OpenStudy (jdoe0001):

\(\bf \cfrac{x^2yz}{y^2}\cdot \cfrac{y}{2x}\implies \cfrac{x^2yz\cdot y}{y^2\cdot 2x}\implies \cfrac{x^2yzy}{y^2x}\implies \cfrac{\cancel{x} x\cancel{y} z\cancel{y}}{\cancel{yy}2\cancel{x}}\implies ?\)

OpenStudy (drewbooks):

xz/2 Thank you very much. :-)

OpenStudy (jdoe0001):

yw